OpenStudy (anonymous):

How to work out, Integral of (3-x)((7^(3-x)^2)) dx?

zepdrix (zepdrix):

\[\Huge \int\limits(3-x)7^{\color{#CC0033}{(3-x)^2}}\;dx\] Let u= this red part.

zepdrix (zepdrix):

\[\Large du=?\]:D

OpenStudy (anonymous):

\[\frac{ -7^{(3-x)^{3}} }{ 3\ln(7) }\]

zepdrix (zepdrix):

Don't simply post answers. Help the person through the problem -_-

OpenStudy (anonymous):

is du, (3-x)^2 and do you bring the 2 down and make if 2(3-x)?

OpenStudy (anonymous):

the suspense was killing me >.<

OpenStudy (anonymous):

yes with dx at the end

OpenStudy (anonymous):

Would u distribute the 2 in?

OpenStudy (anonymous):

and a - in front

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so would dx be du/-2?

OpenStudy (anonymous):

du = -2(3-x)dx

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

remember with u sub you try to substitute all x out

OpenStudy (anonymous):

why do you need a - in the front?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

you also take a derivative of the inside and multiply

OpenStudy (anonymous):

derivative of 3 = 0 + derivative of -x = -1

OpenStudy (anonymous):

go to this website watch some videos under chain rule great explanations. http://patrickjmt.com/

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem, that website is a lifesaver for math